23 Find the critical value SELECT ALL APPLICABLE CHOICES A) 4.075 B) 3.075 C) 31.319 D) 31.653 corresponding to a sample size of 15 and a confidence level of 99.0 percent 24 The confidence Interval for the standard SELECT ALL APPLICABLE CHOICES deviation is given by A) True B) False (n- 1)s2 25 The confidence Interval for variance is given SELECT ALL APPLICABLE CHOICES A) False B) True 26 SELECT ALL APPLICABLE CHOICES A) z - distribution B) t - distribution ơis UNKNOWN n 30 then we use 27
Answer #1

23) 99% confidence interval indicates that the significance level of 0-1-99/100 0.01 .

This is the total area under both right and left tails of chi^2 distribution.

The are under the right tail is 0.01/2=0.005

The degrees of freedom is n-1=15-1=14, for a sample of size n=15

We go to the chi-square table and get a chi-square value of 31.3193 for df=14 for area under the right tail 0.005

This is the right tail critical value 2e

ans: C) 31.319

24) the confidence interval for variance sigma^2 is

(n-1)s(n -1)s 2

and for the standard deviation is

(n-1)s2 (n-1)s2 2 2 XR

Hence the confidence interval given is not for standard deviation.

ans:B) False

25) The confidence interval for variance is

(n-1)s(n -1)s 2

The confidence interval given ar{x}-E<mu<ar{x}+E is for mean

ans: A) False

26) We use the z distribution if sigma is known or n>30.

Since we have n>30, one of the above conditions is met. Hence we can use z-distribution.

ans: A) z-distribution.